OpenStudy (anonymous):

GAIN A MEDAL AND A FAN Part a: write the function rule for each function reflected in the given axis f(x) = x+1 ; x axis f(x) = 3x ; x axis f(x) = 2x - 4 ; x axis Part b: Write an equation for each transformation of y=x Vertical stretch by a factor of 4 vertical stretch by a factor of 2 vertical compression by a factor of 1/2 vertical compression by a factor of 1/4

OpenStudy (anonymous):

to reflect about the x axis, find -f(x).

OpenStudy (anonymous):

\[g(x) = a\cdot f \left(x-h \right)+k\] where a is a vertical stretch if |a|>1 and vertical compression if 0<|a|<1 a reflection about the x axis if a < 0 (i.e. if a is negative) a horizontal shift of h and a vertical shift of k. remember for the horizontal shift, h is the value of x which make (x-h) = 0